Please help me with this problem as I am terrible at math. It comes from a worksheet.
John drove to his uncle’s house in a distant city and the trip took 5 hours. When he returned home, there was less traffic and the trip only took 3 hours. If John drove 26 miles per hour faster on the return trip, how fast did he drive each way?

Answer by ankor@dixie-net.com(15746)

John drove to his uncle’s house in a distant city and the trip took 5 hours. When he returned home, there was less traffic and the trip only took 3 hours. If John drove 26 miles per hour faster on the return trip, how fast did he drive each way?
:
Let s = speed to
Then
(s+26) = speed back
:
We know the distance there and back is the same so we can make a simple equation from that fact:
Distance = time * speed:
:
To dist = back dist
5*s = 3(s+26)
5s = 3s + 78
5s - 3s = 78
2s = 78
s = 78/2
s = 39 mph to, and 39+26 = 65 mph on the return trip
;
:
A good way to check our solution is find the distances, they should be equal:
5 * 39 = 195
3 * 65 = 195 also, proves our solutions
:
That's not so hard, is it? Did it make sense to you?